OpenStudy (anonymous):

If a 36-foot tree casts a 28-foot shadow at the same time near by building casts a 70-foot shadow, how tall is the building?

OpenStudy (anonymous):

hello radar!

OpenStudy (radar):

36/28 = x/70 set up a proportion. 28x=70X36

OpenStudy (radar):

Hey satellite73, whats happening?

OpenStudy (anonymous):

i got 90

OpenStudy (radar):

Hikari93, can you take it from there. If not let me know. One thing we can say, looking at the problem the building is taller than the shadow because the tree is taller than its shadow.

OpenStudy (radar):

I got 90 also.

OpenStudy (anonymous):

\[\theta =atan(36/28)=52.139\]\[70\times \tan \theta =90.046\approx 90\]